If f(4) equals -1 and the graph of f is parallel to the line 3x-8y=16, find the equation for the linear function and write the answer in y=mx+b form.


Thanks so much for your help! My test is tomorrow

Answer by ewatrrr(24785)

 
Hi,
using y = mx = b
the graph of f is parallel to the line 3x-8y=16 OR y = (3/8)x - 2 , m = 3/8
Line f(x) = (3/8)x + b
f(4) = 3/2 + b = -1 or b = -5/2
Line f(x) = (3/8)x - 5/2